OpenStudy (anonymous):

express y=sqrt of x^2+5 as y=f(g(x))

OpenStudy (anonymous):

\[ y=\sqrt{x ^{2}+5}\] as y=f(g(x))

OpenStudy (campbell_st):

let g(x) = x^2 + 5 and \[f(x) = \sqrt{x} \]

OpenStudy (anonymous):

So it would be written how? That's where i'm confused how do we write it

OpenStudy (campbell_st):

\[f(g(x)) = f(x^2 + 5) = \sqrt{x^2 + 5}\]

OpenStudy (anonymous):

okay so do you know how to write y=f(u) and u=g(x) ? Sooo confusing where does u play a part in this?

OpenStudy (campbell_st):

well \[y = \sqrt{u} .... and..... u = x^2 + 5\]

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhh.... okay I get what you did. One quick dumb question haha how do we know that y=\[\sqrt{u}\]

OpenStudy (campbell_st):

its just the most obvious....

OpenStudy (anonymous):

okay I got it!!! Thank you for the help